What is the temperature coeefficient of resistance of a piece of wire having a resistance of 12. 85 ohm at 30 degrees celcius and 6.85 ohm at 0 degrees celcius. what will be its resistance?


1
Expert's answer
2021-05-03T10:34:08-0400
"R_2=R_1(1+\alpha(T_2-T_1))\\12.85=6.85(1+\alpha(30-0))\\
\alpha=0.0292\ K^{-1}"